Circular No. 8349

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q4 (NEAT) J. Lecacheux, LESIA, Meudon Observatory; and E. Frappa, St-Etienne

Planetarium, write: “We have observed the concentric dust shells of comet C/2001 Q4 with the Pic-du-Midi 1.05-m reflector during six 3-hr sessions from May 14 to 19, around the time of perihelion. Direct measurement of the radial expansion on consecutive exposures yields 163 ± 20 m/s. From the measured 12000-km shell interval, a first approximation of P = 20.5 ± 3 hr can be inferred for the period of nucleus rotation, suggesting the exclusion of any value shorter than 17–18 hr. Then by blinking images of the inner shell taken at 1-day or several-day interval(s), and assuming constant expansion velocity, we obtain a refined probable period P = 23.2 ± 0.25 hr. We also followed a small dust jet rotating counterclockwise (facing the sun) on May 14; its estimated angular motion of ∼ 16 deg/hr agreed with the above period. Fibrous-like features, not more 300 km wide, appear within the brightest shell at 10000–15000 km from the nucleus; they show a pure radial expansion and no other morphology change in 3 hr, and they recur in the following shell 0.9 day later. In fact, we observe a complex of 3 or or 4 muddled components (sub-shells), issued from so many active regions and with slightly different expansion rates.”

Further visual total-magnitude estimates: May 19.87 UT, 4.2 (V. Znojil, Brno, Czech Republic, naked eye); 22.91, 4.5 (M. Meyer, Ursenbach, Germany, naked eye); 24.91, 4.7 (A. Diepvens, Balen, Belgium, 20×50 binoculars); 27.90, 5.0 (A. Amorim, Florianopolis, Brazil, 7×50 binoculars); 30.92, 5.4 (J. J. Gonzalez, Asturias, Spain, 7×50 binoculars).

SUPERNOVA 2004by IN NGC 7116 M. Armstrong, Rolvenden, U.K., reports his discovery of an appar-

ent supernova (mag 18.9) on unfiltered CCD images taken on May 24.102 and 28.026 UT with a 0.35-m reflector. SN 2004by is located at α = 21^{h}42^{m}41^{s}.88, δ = +28^{o}56′50′′.6 (equinox 2000.0), which is ≈ 16′′ east and 1′′.9 north of the center of NGC 7116. Nothing was visible at this position on his images taken on 2001 July 15 and 2003 Dec. 17 (limiting mag ∼ 19.5) and on Palomar Sky survey red and blue plates. M. Mobberley, Cockfield, U.K., confirmed the new object on an unfiltered CCD image taken on May 27.951 with a 0.35-m reflector.

2004 May 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green